OpenStudy (anonymous):

Stats! How to know if this binomial distribution is approximately normal. n=20 P=.2

OpenStudy (anonymous):

But how do I know its distributed approximately normal.

jimthompson5910 (jim_thompson5910):

If both np > 5 and n(1-p) > 5, then it's approximately normal

jimthompson5910 (jim_thompson5910):

some books use a more strict guideline: np > 10 and n(1-p) > 10

OpenStudy (anonymous):

Oh I recall learning that in class, thank you Jim_thompson. Also I noticed that I can't fan you

jimthompson5910 (jim_thompson5910):

oh for some reason i blocked you in the past, i forget why. You're unblocked now though

OpenStudy (anonymous):

lol XD

OpenStudy (anonymous):

I don't really remember I did something wrong

jimthompson5910 (jim_thompson5910):

I forget so I have no clue. I usually block if someone is harassing me or something

jimthompson5910 (jim_thompson5910):

but np = 20*0.2 = 4 which is NOT greater than 5. So np > 5 is false meaning that the binomial distribution is not approximately normal

jimthompson5910 (jim_thompson5910):

to fix this, you would need to increase n

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

It doesn't actually tell me if I need to fix it, it's just a yes or no question

jimthompson5910 (jim_thompson5910):

so it's just "no" because np = 20*0.2 = 4 which means np > 5 is false. You should check n(1-p) as well, but if one is false, then the whole thing doesn't work.

OpenStudy (anonymous):

alright